OpenStudy (anonymous):

calculate the following limits and represents the result you get (represents the asymptotes and associated curves in appropriate cases). \[\lim_{x \rightarrow + \infty} \left( x^{2}/3 - 2x \right)\]

OpenStudy (anonymous):

is this \[\frac{x^2}{3-2x}\]?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

\[\frac{x^2}{3}-2x\]?

OpenStudy (anonymous):

\[\frac{ x2 }{ 3 } - 2x\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

this is a quadratic polynomial, therefore the graph is a parabola that opens up (because the leading coefficient is positive) therefore as \(x\) gets larger, it goes to infinity as well

OpenStudy (anonymous):

i dont understand..

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=y%3Dx^2%2F3-2x

OpenStudy (anonymous):

as \(x\) gets larger, so does \(\frac{x^2}{3}-2x\) try it with \(x=10,000\) and see what you get

OpenStudy (anonymous):

i got -33326